OpenStudy (anonymous):

Y/5 + 2 = -8 ? Show steps please

OpenStudy (anonymous):

Are you solving for y? If so, the answer is -50. Step 1: Plug -50 into Y -50/5 + 2 = -8 Step 2: Solve according to PEMDAS -50/5 + 2 = -8 -10 + 2 = -8 -8 = -8